Solve the following differential equations. Be sure you can interpret them in terms of periodic motion, and graph the solution.

3.    x′′+2x′+x= t + δ(t);    x(0)=0,x′(0)=1
Answer #1

X+2x+x=t+sal - taking caplace from form on bom side LX (8)8-Xo) 8 - Xo) 12X(s) - 2Xo) tX(4) = $2 + 1 Apply given condition X